OpenStudy (anonymous):

the no of positive integral solution of abc=30 A) 30 B)27 C)8 D)none of these

OpenStudy (anonymous):

1x1x30;2x3x5;10x3x1;2x15x1;5x6x1 therefore, total cases= 5x3=15

OpenStudy (anonymous):

no the the result is another the result is 27 but i don't know how i get the answer

OpenStudy (anonymous):

30 = 2*3*5 => 3! = 6 ways 30 = 1*1*30 => 3!/2! = 3 ways 30 = with just 1's => 3 *3c1*3 = 18 ways total = 6+3+18 = 27 ways

OpenStudy (anonymous):

Strange question, I would not have said myself that 2*3*5 and 2*5*3 were different solutions....